An investment of ​$130 comma 000 was made by a business club. The investment was split into three parts and lasted for one year. The first part of the investment earned​ 8% interest, the second​ 6%, and the third​ 9%. Total interest from the investments was $ 9540. The interest from the first investment was 2 times the interest from the second. Find the amounts of the three parts of the investment.
Answer by Theo(13342)

total investment is $130,000.
investment is split into 3 parts.
first part earns 8%.
second part earns 6%
third part earns 9%.
total interest from the investments = $9540.
interest from first part = 2 * interest from second part.

let a = the amount invested in the first part.
let b = the amount invested in the second part.
let c = the amount invested in the third part.

you have two equations that need to be solved simultaneously.

they are:

a + b + c = 130,000
.08 * a + .06 * b + .09 * c = 9540

you are given that interest from the first investment is 2 times the interest from the second investment.

equation for that is .08 * a = 2 * .06 * b
divide both sides of that equaiton by .08 to get:
a = (2 * .06 * b) / .08
simplify that equation to get a = .12 * b / .08.
simplify further to get a = 1.5 * b

in the equation of a + b + c = 130,000, replace a with 1.5 * b to get a + b + c = 130,000 which becomes 1.5 * b + b + c = 130,000 which becomes 2.5 * b + c = 130,000

the first equation has now becomes 2.5 * b + c = 130,000

replace a with 1.5 * b in the second equation to get .08 * a + .06 * b + .09 * c = 9540 becomes .08 * 1.5 * b + .06 * b + .09 * c = 9540 which becomes .12 * b + .06 * b + .09 * c = 9540 which becomes .18 * b + .09 * c = 9540.


your 2 equations to solve simultaneously are now:

2.5 * b + c = 130,000
.18 * b + .09 * c = 9540

multiply both sides of the first equation by .09 and leave the second equation as is to get:

.225 * b + 09 * c = 11700
.18 * b + .09 * c = 9540

subtract the second equation from the first to get:

.045 * b = 2160

solve for b to get b = 2160 / .045 = 48,000

since a = 1.5 * b, then a = 72,000

that makes c = 10,000 because 130,000 - 48,000 - 72,000 = 10,000

a + b + c becomes 72,000 + 48,000 + 10,000 = 130,000, which is true.

.08 * a + .06 * b + .09 * c = 9540 becomes .08 * 72,000 + .06 * 48,000 + .09 * 10,000 = 9540 which becomes 5760 + 2880 + 900 = 9540 which becomes 9540 = 9540, which is true.

your solution is that the amount of the three parts of the investments are 72,000 at 8%, 48,000 at 6%, 10,000 at 9%.